OpenStudy (pulsified333):

Find the derivative equation of: y=−2.5x^2+0.5x−1.42 I thought 5x+(1/2) was the answer but it is not

OpenStudy (pulsified333):

@dan815 @Zarkon

OpenStudy (pulsified333):

@mathstudent55

OpenStudy (michele_laino):

hint: we have these subsequent derivatives: \[\Large \begin{gathered} {x^2} \to 2x \hfill \\ \hfill \\ x \to 1 \hfill \\ \end{gathered} \]

OpenStudy (pulsified333):

subsequent?

OpenStudy (michele_laino):

yes! I wrote the derivatives of the functions \(x^2\) and \(x\) respectively

OpenStudy (michele_laino):

hint: the derivative of the first term is \(-2.5 \cdot 2 \cdot x\)

OpenStudy (pulsified333):

oh

OpenStudy (pulsified333):

so its just -5x+1/2